OpenStudy (anonymous):

How do you use the second fundamental theorem of calculus to find a derivative of the function of an integral from -4 to sin(x)?

OpenStudy (anonymous):

\[\int\limits_{-4}^{\sin(x)}[\cos(t^5)+t] dt\]

OpenStudy (amistre64):

not sure what the second one is, but i think you are looking for: \[\fracd{dx}\int_{a}^{b}f(t)~dt=f(b)b'-f(a)a'\]

OpenStudy (anonymous):

The second part is the actual question that I have. I'll write out what the second fundamental theorem says

OpenStudy (anonymous):

if \[f(x)=\int\limits_{a}^{x}g(t)dt \] then \[f'(x)= g(x)\]

OpenStudy (anonymous):

I can find the derivative with respect to sin(x) but not just x. I think. What I got was \[f'(\sin(x))= \frac{ \cos(\sin^5(x))+\sin(x) }{ \cos(x) }\]

OpenStudy (amistre64):

\[\frac d{dx}\int\limits_{-4}^{\sin(x)}[\cos(t^5)+t] dt\]\[\hspace{5em}=[~cos(sin^5(x)+sin(x)~]~cos(x) dt\]

OpenStudy (anonymous):

why is it multiplied instead of divided?

OpenStudy (anonymous):

by cos(x)

OpenStudy (amistre64):

chain rule

OpenStudy (anonymous):

Yes, but going back a step, wouldn't it be \[f'(sinx)*\cos(x)=\frac{ d }{ dx }(\int\limits_{-4}^{\sin(x)}(\cos(t^5)+t) dt)\]?

OpenStudy (amistre64):

by the definition of integrating \[\int_{a(x)}^{b(x)} f(t)~dt=F(b(x))-F(a(x))\] and by the definition of derivatives \[D_x[F(b(x))-F(a(x))]\] \[=D_x[F(b(x))]-D_x[F(a(x))]\] \[=f(b(x))*b'(x)-f(a(x))*a'(x)\]

OpenStudy (amistre64):

your a'=0 so that goes away

OpenStudy (amistre64):

so you are left with f(b)*b'

OpenStudy (amistre64):

b=sinx; b'=cosx

OpenStudy (amistre64):

i think you are trying to make it more difficult simply because f' is not very pretty

OpenStudy (anonymous):

Ive never seen the definition of a derivative written like that before. I'm having a hard time interpreting. But I still don't understand why the cos(x) is being multiplied instead of divided, and if that is f'(x)? I thought it was f'(sin(x))

OpenStudy (amistre64):

does the chain rule pop out a division or a multiplication?

OpenStudy (anonymous):

Multiplication, but when it changes sides of the equation, it'd be division

OpenStudy (amistre64):

why would it change sides of the equation?

OpenStudy (anonymous):

To get f'(x) by itself

OpenStudy (anonymous):

I'm having a hard time with this concept because I was only taught through a video due to a strike. The textbook isn't making it easier.

OpenStudy (amistre64):

your confusing something .... show me your steps so i can see what square peg your trying to cram into the round hole

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Since the upper bound is sin(x), I'm getting f(sin(x)) instead of f(x). That's my first issue. I have it written like this: \[f'(\sin(x))*\cos(x)=(\cos(\sin^5(x))+\sin(x))\]

OpenStudy (anonymous):

Which is why I'm confused when it comes to the cos(x) being multiplied on the other side like you have it

OpenStudy (amistre64):

would you agree that:\[F(x)=F(b)-F(a)\]

OpenStudy (amistre64):

\[F(x)=F(sin(x))-F(-4)\]

OpenStudy (anonymous):

I'm supposed to be using the second fundamental theorem, which I stated near the top of this... THats the first fundamental theorem

OpenStudy (amistre64):

there is no appreciable difference between the two of them

OpenStudy (amistre64):

the second is simply stating the consequences of the first if the lower limit is a constant

OpenStudy (anonymous):

I can vaguely see that... Let me process for a minute.

OpenStudy (anonymous):

Okay can you help me solve it using that and then help me translate it to the way that it is written for the Second Fundamental Theorem?

OpenStudy (amistre64):

sure, in fact im using the the first to establish that F(x) is, then by deriving F(x) we get to the second thrms by default

OpenStudy (anonymous):

so now we just differentiate both sides?

OpenStudy (anonymous):

Oh wait.... the derivative.... what... F'(x)=F'(sin(x))??

OpenStudy (anonymous):

Oh chain rule

OpenStudy (anonymous):

so F is the integral function, right? And so F'(x)=F'(sinx)*cos(x)?

OpenStudy (amistre64):

yes, \[\Large F(x)=\int_{k}^{sinx}f(t)~dt\] \[\Large F(x)=F(sin(x))-F(k)\] \[\Large F'(x)=F'(sin(x))(sinx)'-F'(k)k'\] since k', the derivative of any constant, is 0 \[\Large F'(x)=F'(sin(x))(sinx)'\] \[\Large F'(x)=F'(sin(x))(cosx)\] and F' = f

OpenStudy (amistre64):

therefore\[F'(x)=f(sinx)~cosx\]

OpenStudy (anonymous):

But f(x) is defined by an integral... so I can't type that in... because\[f(x)= \int\limits_{-4}^{sinx}(\cos(t^5)+t)dt\] right?

OpenStudy (anonymous):

oh wait no...

OpenStudy (amistre64):

no ...

OpenStudy (anonymous):

f(x) is the inside of the integral, right?

OpenStudy (amistre64):

f(t) is the inside of the integral; and the integral itself become F(x) the conventional use of upper and lower case f might be causing you some grief

OpenStudy (anonymous):

Okay that is the right answer... But I am still confused how to approach it from the other direction

OpenStudy (amistre64):

the derivative of F is f, and we are looking for the derivative of F

OpenStudy (anonymous):

I understand the way that we just did it now.

OpenStudy (amistre64):

if i were to try to adapt your setup \[f(x)=\int\limits_{a}^{x}g(t)dt\] \[f'(x)=g(x)\] thats fine, but lets define the limit as a function of x, instead of just x; say u(x) = x .... then we have to play with a chain rule \[f(x)=\int\limits_{a}^{u(x)}g(t)dt\] \[f'(x)=g(u(x))*u'(x)\]

OpenStudy (amistre64):

doesnt that revert to: f'(x) = g(x)? since u(x) = x ; and u' = 1

OpenStudy (anonymous):

Man, that's ugly.

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

You've been a huge help. But, the way that mine is set up, why isn't it ending up with f'(u(x))?

OpenStudy (amistre64):

becasue we are defining f'(x) in terms of g(u(x)) and u'(x) your setup seems a little better at expressing the point

OpenStudy (amistre64):

if we ignore the (x) parts we can prolly read this even better f' = g(u) * u'

OpenStudy (anonymous):

Umm nope, it was making more sense with the 'x's to me.

OpenStudy (amistre64):

lol, then leave them in ;)

OpenStudy (anonymous):

Haha but even though it makes *more* sense, it doesn't quite click.

OpenStudy (anonymous):

Because to me, you have x's in one equation, then you change one x to u(x), why doesnt the other x also change to u(x)?

OpenStudy (amistre64):

no, go thru the long way, integrate up, insert the limits, and derive it back down .... that is the best way to understand it.

OpenStudy (anonymous):

Argh. I'll ask my prof to help me through in more detail tomorrow. Thank you! Would you mind helping with one that is a definite integral?

OpenStudy (amistre64):

\[f(x)=\int\limits_{a}^{u(x)}g(t)dt\] \[f(x)=G(u(x))-G(k)\] \[f'(x)=g(u(x))~u'(x)-g(k)\cancel{k'}^0\] \[f'(x)=g(u(x))~u'(x)\]

OpenStudy (amistre64):

ill help

OpenStudy (amistre64):

in the first one, we are not replacing x with u(x), its just a consequence of the integration is all

OpenStudy (amistre64):

i mean f(x) is not equal to f(u(x))

OpenStudy (anonymous):

I think I need to think about that one by myself for a while :P Okay. I already know the answer, I just cant get to it with these concepts. \[\int\limits_{0}^{1}t^2dt\]

OpenStudy (amistre64):

this ones pretty basic, what do you need help with?

OpenStudy (anonymous):

I know that it's basic, I want to understand it so I can apply the same steps to a harder one. I'm supposed to use the set up that I gave you to evaluate it.

OpenStudy (anonymous):

Which means I need x's on the boundaries, right? So I can split it into 2 integrals

OpenStudy (amistre64):

you dont need xs ... we you trying to determine the derivative of the inegration again? or are you just trying to find the value of the integration itself?

OpenStudy (anonymous):

Just the value. Sorry.

OpenStudy (amistre64):

do you recall the power rule for derivatives?

OpenStudy (anonymous):

And not using integration techniques. I mean I know how to integrate that.

OpenStudy (amistre64):

not using integration techniques .... then what shall we try to use?

OpenStudy (amistre64):

the limit of reimann sums?

OpenStudy (anonymous):

The second fundamental theorem of calculus.

OpenStudy (anonymous):

We've used riemann sums to prove this one in class.

OpenStudy (anonymous):

I know it's sketchy going to a site introduced to you by a stranger on the internet, but cosketch is nice and clean and easier than typing these things out. It's just drawing... Could we use that?

OpenStudy (amistre64):

the 2nd fundamental thrm states: \[if~~f(x)=\int\limits_{a}^{x}g(t)~dt~~then~~f'(x)=g(x) \] im not sure how we would use the second thrm to determine the value of the definite integration without actually intgrating it

OpenStudy (anonymous):

Neither am I, but we haven't learned the reverse chain rule and the original problem I have would need that to integrate. So I know it can be done.

OpenStudy (anonymous):

That's why I said that we needed x on the boundary

OpenStudy (anonymous):

Oh! I misread! It says the First! I think I can solve it now.

OpenStudy (amistre64):

in any case, you are going to need to integrate it up to a function that derives down to t^2

OpenStudy (amistre64):

power functions derive down a degree and integrate up a degree so there is some function k t^3 that has a derivate t^2 k t^3 derives to 3k t^2; what does k have to be in order for this to equal t^2?

OpenStudy (anonymous):

Yeah I dont know how to solve without integrating.... I Know how to integrate it. But I'm not supposed to.

OpenStudy (anonymous):

one third.

OpenStudy (amistre64):

i think you need to integrate it, but then apply the first thrm to find the value

OpenStudy (anonymous):

But the question that I have to solve is not integrable with the rules that we have learned. And we should be able to solve it.

OpenStudy (amistre64):

so, F(t) = 1/3 t^3; and the first thrm says that the value is equal to F(b)-F(a)

OpenStudy (anonymous):

I have a second, messy one. I just wanted to see if I could get the answer through these methods for an easy one first.

OpenStudy (anonymous):

We arent supposed to integrate.

OpenStudy (amistre64):

youre going to have to clear that up with your teacher than.

OpenStudy (anonymous):

Okay. -sigh- Thanks anyways.

OpenStudy (amistre64):

good luck with it :)